calculate the amount of solid (in grams) that is produced when 18.00 mL of 0.1100 M sodiun chromate is mixed with 21.00 mL of 0.09300 M aluminum chloride.

write the balanced molecular and net ionic equations for this reaction. State the spectator ions left in solution
Answer #1

Balanced molecular equation: 2AlCl3(aq) + 3Na2CrO4(aq) = Al2(CrO4)3(s) + 6NaCl(aq)

Net ionic equation: 2Al3+(aq) + 3CrO42-(aq) = Al2(CrO4)3(s)

2 Aliz (aq) + 8 Naz Go On (aq) = Al₂ C Grou) (8) + 6 Nallfar) C = 0.093M C = O. JIM = 21 ml = 0.024 V = 18 ml > 0.0212 = 0.01... No. of moles of h Alz (Co Ou), produced from Naz Gron = /1.98% 10 3 x 1 mol = 0.66%d0-3 anol. This I comes from the pit-rSpectator ions left in the solution are Na+ and Cl-.